OpenStudy (anonymous):

Will medal and fan!! Order the numbers from least to greatest: -7/8, 7/4, -15/2, and -13/16

OpenStudy (anonymous):

@KendrickLamar2014

OpenStudy (anonymous):

-13/6, -7/8, 7/4, 15/2

OpenStudy (anonymous):

its 13/16

OpenStudy (anonymous):

-13/16, -7/8, 7/4, 15/2

OpenStudy (anonymous):

are u sure -13/16 is less than -15/2

OpenStudy (anonymous):

lol i know 7/4 is the largest cuz its not negative and i think -15/2 is the least cuz its the only negative mixed number i'm just not sure about the middle

OpenStudy (anonymous):

oh sorry. i didnt notice it was -15/2

OpenStudy (anonymous):

its ok

OpenStudy (anonymous):

so it would be -13/16, -15/2, -7/8, 7/4

OpenStudy (tkhunny):

-7/8 = -14/16 7/4 = 28/16 -15/2 = -120/16 -13/16 = -13/16 It's a LOT easier, like that.

OpenStudy (anonymous):

so -15/2, -7/8, -13/16, 7/4 right?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

ok thanks :)

OpenStudy (anonymous):

i like ur profile pic btw

OpenStudy (anonymous):

yw :)

OpenStudy (anonymous):

sorry i gave you the wrong answer at first

OpenStudy (anonymous):

its ok

OpenStudy (tkhunny):

Note: It's not always this easy to get them all to the same denominator. This problem was exceptionally convenient.